### v1

@@ A.1 Benchmarking on UCI dataset
In this section, we compare our method to Ensemble and Evidential using the UCI benchmark dataset. This is intended to provide readers with a direct comparison to Lakshminarayanan et al., (2017) and Amini et al., (2020) on the same dataset used in both works. The collection consists of nine real world regression problems, each with 10–20 features and hundreds to tens of thousands of observations. We note the Wine dataset within UCI contains discrete values (ratings of wine characteristics, such as color and taste) which may render the assumption of a continuous, symmetrical data distribution less appropriate if these values are skewed. More recently, larger UQ datasets have been published in Dheur and Ben Taieb, (2023), which may be useful in assessing the state-of-the-art in non-time-series UQ methods. We follow Lakshminarayanan et al., (2017) and Amini et al., (2020) in evaluating our method on root mean squared error (RMSE, which assesses forecast accuracy) and NLL (which assesses overall distributional fit), and compare against Ensemble and Evidential. While we do not explicitly compare inference speed, as our Combined method also uses ensembling, inference speed is expected to be comparable to Ensemble while being slower than Evidential. We use the source code provided by Amini et al., (2020), with the default topology of a single hidden layer with 50 units for both Ensemble and Evidential [^11] Source code for Amini et al., (2020) is available on Github: https://github.com/aamini/evidential-deep-learning. For Combined, as individual modelling of distribution parameters (Section 3.2) requires a network with two or more hidden layers, we have used a single hidden layer with 24 units, followed by 4 separate stacks of a single hidden layer with 6 units each. Thus, the total number of non-linear units is 48 (compared to 50 for Ensemble and Evidential). Note that even though the total number of units are similar across the three models, learning capacity may differ due to different topologies.
@@
Table 4 records experiment results on the UCI dataset. On RMSE, we find that both Ensemble and Combined have performed well, having the best RMSE in four datasets each. In two of the sets (Kin8nm and Naval), all three methods produced highly accurate results that are not separable to two decimal points. Turning to NLL, we observe a trend towards Combined having lower NLL than the other two methods for four sets, followed by Ensemble with three sets. Comparing Combined to Evidential, we find that Combined generally has lower RMSE (7 of 9 sets) and NLL (6 of 9 sets). Although our method is designed for UQ of complex time-series and all 9 datasets are pooled (non-time-series) datasets, we still observe some improvements in both RMSE and NLL.
@@
Next, we present further ablation studies on the UCI dataset. Table 5 records results of Alternative, which utilizes ensembling and SMD parameterization but not separate modelling of hyperparameters. Alternative has the same network topology as Ensemble and Evidential (a single hidden layer with 50 units), as opposed to Combined which has two hidden layers with a total of 48 units. We observe that Ensemble has the lowest RMSE in 5 (of 9) datasets, followed by Alternative (3 of 9), while Alternative has the best NLL in 6 (of 9) datasets and Ensemble has 3 (of 9). On both metrics, Evidential has the least favourable performance. Comparing Combined in Table 4 and Alternative in Table 5, Combined has lower RMSE and NLL in 5 of 9 datasets. Thus, we conclude that separate modelling of hyperparameters provided an incremental benefit on the UCI datasets.
@@
Table 5: Comparing Ensemble, Evidential and Alternative (without separate modelling of the four parameters of SMD) on RMSE and NLL using the UCI benchmark datasets. Average result and standard deviation over 5 trials for each method. The best method for each dataset and metric is highlighted in bold.
@@
Dataset Ensemble Evidential Alternative Ensemble Evidential Alternative
@@
In Table 6, we further remove model averaging. The network used is identical to Evidential but trained using the SMD parameterization (i.e., we simply change the loss function in Evidential to Equation (12)). We observe that the network trained using the SMD parameterization has lower RMSE in 6 of 9 and lower NLL in 8 out 9 datasets. We argue that the improved performance of the SMD parameterization is due to its simplicity.
Table 6: Comparing Normal-Inverse-Gamma and Normal-Gamma on RMSE and NLL using the UCI benchmark datasets. Average result and standard deviation over 5 trials for each method. The best method for each dataset and loss function is highlighted in bold.
@@
Appendix B Further analysis of parameters in a Scale Mixture
In the network architecture proposed in Section 3, output of the network is $\bm{\zeta}=(\gamma,\sigma^{2},\alpha,\beta)$, which parameterises the SMD (Equation (10)). However, as noted in Section 3, we can set $\alpha=\beta$ and reduce the number of parameters to three (Equation (14)). Thus, an alternative specification of the network is to output $\bm{\zeta}=(\gamma,\sigma^{2},\alpha)$ (i.e., three parameters instead of four and are computed through three subnetworks, instead of four in Figure 1). We label this network A=B. In Table 7, we compare Combined (4 parameters) with S2B (3 parameters) using the UCI dataset (as introduced in Section A.1). We observe that A=B is better than Combined on 8 (of 9) datasets on RMSE, while Combined is better than A=B on 1 (of 9). On NLL, A=B is better than Combined on 5 (of 9) datasets, while Combined is better than A=B on 4 (of 9). Even though A=B has a higher number of datasets with lower RMSE and NLL, we note that the differences are very small and are within margin of error (due to randomness in neural network training). Thus, we conclude that the two methods provide near identical results but note that A=B is simpler and more interpretable. However, we choose Combined with four subnetworks to conduct our analysis so that parameters can also be compared with those from Evidential.

### v2

@@ B.1 Comparison in UCI datasets
We follow Lakshminarayanan et al., (2017) and Amini et al., (2020) in evaluating our method using RMSE which assesses forecast accuracy and NLL which assesses overall distributional fit, and compare the measures against Ensemble and Evidential. While we do not explicitly compare inference speed, as our Combined method also uses ensembling, inference speed is expected to be comparable to Ensemble while being slower than Evidential. We use the source code provided by Amini et al., (2020), with the default topology of a single hidden layer with 50 units for both Ensemble and Evidential [^16] Source code for Amini et al., (2020) is available on Github: https://github.com/aamini/evidential-deep-learning. As individual modelling of distribution parameters (Section 2.4.3) in Combined requires a network with two or more hidden layers, we use a single hidden layer with 24 units, followed by 4 separate stacks of a single hidden layer with 6 units each. Thus, the total number of non-linear units is 48 (compared to 50 for Ensemble and Evidential). Note that even though the total number of units are similar across the three models, learning capacity may differ due to different topologies.
@@
Table 4 records experiment results. On RMSE, we find that both Ensemble and Combined have performed well, having the best RMSE in four datasets each. In two of the sets (Kin8nm and Naval), all three methods produced highly accurate results that are not separable to two decimal points. Turning to NLL, we observe a trend towards Combined having lower NLL than the other two methods for four sets, followed by Ensemble with three sets. Comparing Combined to Evidential, we find that Combined generally has lower RMSE (7 of 9 sets) and NLL (6 of 9 sets). Although our method is designed for UQ of complex time-series and all 9 datasets are non-time-series datasets, we still observe some improvements in both RMSE and NLL.
B.2 Ablation studies
@@
We further present ablation studies on the UCI dataset. The first study compares Ensemble and Evidential with Single Output (as in Table 3), which utilizes model averaging and SMD parameterization but not separate modelling of hyperparameters in Combined. Single Output has the same network topology as Ensemble and Evidential (a single hidden layer with 50 units), as opposed to Combined which has two hidden layers with a total of 48 units. We observe from Table 5 that Ensemble has the lowest RMSE in 5 (of 9) datasets, followed by Single Output (3 of 9), while Single Output has the best NLL in 6 (of 9) datasets and Ensemble has 3 (of 9). On both metrics, Evidential has the least favorable performance. Comparing Combined in Table 4 and Single Output in Table 5, Combined has lower RMSE and NLL in 5 of 9 datasets. Thus, we conclude that separate modelling of hyperparameters provided an incremental benefit on the UCI datasets.
@@
Table 5: Comparing Ensemble, Evidential and Single Output (Combined but without separate modelling of the four parameters of SMD) on RMSE and NLL using the UCI benchmark datasets. Average result and standard deviation over 5 trials for each method. The best method for each dataset and metric is highlighted in bold.
@@
Dataset Ensemble Evidential Single Output Ensemble Evidential Single Output
@@
In the second study, we further remove model averaging. The network used is identical to Evidential but trained using the SMD parameterization (i.e., we simply change the loss function in Evidential to Equation (15)). We observe from Table 6 that the network trained using the SMD parameterization has lower RMSE in 6 of 9 and lower NLL in 8 out 9 datasets. We argue that the improved performance of the SMD parameterization is due to its simplicity.
Table 6: Comparing Normal-Inverse-Gamma and Scale Mixture Distribution on RMSE and NLL using the UCI benchmark datasets. Average result and standard deviation over 5 trials for each method. The best method for each dataset and loss function is highlighted in bold.
@@
Appendix C Further analysis of parameters in a Scale Mixture Distribution
In the network architecture proposed in Section 2.4.3, output of the network is $\bm{\zeta}=(\gamma,\sigma^{2},\alpha,\beta)$, which parameterises the SMD (Equation (13)). However, as noted in Section 2.4.2, we can set $\alpha=\beta$ and reduce the number of parameters to three (Equation (18)). Thus, an alternative specification of the network is to output $\bm{\zeta}=(\gamma,\sigma^{2},\alpha)$ (i.e., three parameters instead of four and are computed through three subnetworks, instead of four in Figure 2). We label this network A=B. In Table 7, we compare Combined (4 parameters) with A=B (3 parameters) using the UCI dataset (as introduced in Section B). We observe that A=B is better than Combined on 8 (of 9) datasets on RMSE, while Combined is better than A=B on 1 (of 9). On NLL, A=B is better than Combined on 5 (of 9) datasets, while Combined is better than A=B on 4 (of 9). Even though A=B has a higher number of datasets with lower RMSE and NLL, we note that the differences are very small and are within margin of error (due to randomness in neural network training). Thus, we conclude that the two methods provide near identical results but note that A=B is simpler and more interpretable. However, we choose Combined with four subnetworks to conduct our analysis so that parameters can also be compared with those from Evidential.
@@ Appendix C Further analysis of parameters in a Scale Mixture Distribution
Appendix D Hyperparameters used in Section 3.1.3
In this section, we provide a list of hyperparameter search ranges and mean hyperparameters used to train the neural networks in Section 3.1.1 and 3.1.2. Hyperparameter search was performed on all combinations of hyperparameters. Both cryptocurrency and U.S. equities datasets share the same hyperparameter ranges but with hyperparameter search performed separately.
Table 8: Hyperparameter ranges used in Section 3.1.1 and 3.1.2. The ‘LSTM layers’ hyperparameter is a list, with the length of the list indicating how many LSTM layers were used and each element of the list indicating the number of units of each LSTM layer. Similarly, ‘Hidden layers’ indicate the number of fully connected hidden layers. Each element of the list indicate the dimension of that hidden layer. ADAM is the optimiser proposed by Kingma and Ba, (2015).
Parameter Search range
LSTM layers $\{[16,8],[32,16,8],[32,16],[64,32,16]\}$
Hidden layers $\{[8],[16,8]\}$
Dropout rate $\{0.2,0.3,0.4\}$
Activation ReLU
Batch size 1,000
Batch normalisation Yes
Early stopping Patience 5 / Tolerance 0.0001
Learning rate $\eta$ $0.01$
Optimiser ADAM